# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v3

@@ 6 Conclusion
In addition to discussing the identification of the shocks, we discussed the problem of labelling them, and how to utilize blended identification (cf. Carriero et al., 2024) to solve the problem of weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices. We discussed estimation of the model parameters by the method of penalized maximum likelihood, and proposed a three-phase estimation procedure by adapting the procedure of Lanne et al. (2017) to nonlinear SVAR models. By making use of the results of Saikkonen (2008), we also presented a sufficient condition for establishing stationarity, ergodicity, and mixing properties of our structural STVAR model. While useful for obtaining a well-behaving estimator, stationarity is not, however, required for the applicability of our identification results. The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
Our empirical application studied the macroeconomic effects of the climate policy uncertainty shock and considered monthly U.S. data from 1987:4 to 2024:12. As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics. We fitted a two-regime logistic STVAR model to the data using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable. We found that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU. Our results are, therefore, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
@@ Appendix A Stationarity and ergodicity
Establishing stationarity and ergodicity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties for the estimator, such as consistency and a conventional limiting distribution. Saikkonen (2008) derives a sufficient condition for ergodicity and stationarity of a smooth transition vector error correction (STVEC) model. Kheifets and Saikkonen (2020) make use of the results of Saikkonen (2008) to show that the stationarity condition readily applies for a STVAR model that can be obtained as a special case of the STVEC model of Saikkonen (2008). However, in both Saikkonen (2008) and Kheifets and Saikkonen (2020), the conditional covariance matrix of the process is defined as a weighted sum of positive definite error term covariance matrices of the regimes. As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model cannot be directly obtained from the model of Saikkonen (2008). The applicability of the results of Saikkonen (2008) also requires that the weights are not exogenous but functions of the preceding $p$ observations (although the time $t$ weights are allowed to depend on the time $t$ observation of an IID process independent of $e_{t}$, we do not discuss this option further, as it violates our assumption of $\mathcal{F}_{t-1}$-measurable or nonrandom weights). Nonetheless, we will show in this section that the stationarity condition applies to our model as well.
@@
The following assumption collects sufficient conditions for establishing ergodicity, stationarity, and mixing properties of our structural STVAR model.
@@
Condition (I) states that the JSR of the companion form AR matrices of the regimes is strictly less than one, and it is analogous to Condition (10) of Saikkonen (2008). Note that this condition is sufficient but not (necessarily) necessary for ergodic stationarity of the process. Condition (II) is analogous Assumption 1 of Saikkonen (2008), and it is innocuous in practice but rules out bounded error distributions, for example. Finally, Condition (III) ensures that the assumptions for the transition function imposed by Saikkonen (2008) are satisfied. This condition rules out nonrandom exogenous weights, which are perfectly fine in terms of our main identification results discussed in Section 3.
@@
Theorem A.1 is proven in Appendix A.1. A notable difference of Theorem A.1 to Theorem 1 of Saikkonen (2008) (and Theorem 1 of Kheifets and Saikkonen, 2020) is that our results holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, while the result of Saikkonen (2008) holds everywhere. To be more specific, the result of Theorem A.1 holds with such values of $B_{1},...,B_{M}$ that the resulting impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$, and Lemma 1 shows that this holds everywhere expect in regions of $\mathbb{R}^{d\times dM}$ that have a zero Lebesgue measure.
Assumption A.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption A.1(I) is computationally demanding to verify in practice (e.g., Chang and Blondel, 2013). This makes it poorly suitable for restricting the parameter space in the numerical maximum likelihood estimation of parameters discussed in Section 4. During estimation, this condition would have to be verified repeatedly a very large of times, thus, substantially increasing the computational cost of the estimation, making it extremely tedious. Therefore, it is useful make use of the following condition in the estimation, which is necessary for Assumption A.1(I).
@@
Condition A.1 states that the usual stability condition is satisfied by each of the regimes. It is necessary for Assumption A.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein). Since our estimation method discussed in Section 4 can produce a number of local solutions, the researcher can choose the best appropriate local solution whose stationarity can be verified by Assumption A.1(I).
Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), which works well for small matrices, but can be computationally extremely demanding if the matrices are large and a tight bound is required. The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
@@ B.1 Proof of Lemma 1
Before proving Lemma 1, we establish the following lemma that made use of in the proof:
@@
Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variable $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
@@
To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$. First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$. Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible. The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$. Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$. Denote this polynomial as $P_{t}(B_{1},...,B_{M})\equiv\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})$. Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero. It then follows from Lemma B.1 that $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
@@ B.2 Proof of Lemma 2
The first part of our proof is largely similar to the proof of Proposition 1 in Lanne et al. (2017) making use of results from independent component analysis (Comon, 1994). First, note that from Equation (3.2), we can rewrite the model as
$y_{t}-\phi_{t}-\boldsymbol{A}_{t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (B.2)
where $\phi_{t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ $(d\times 1)$, $\boldsymbol{A}_{t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{t}$ and $\tilde{\boldsymbol{A}}_{t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{t}-\tilde{\boldsymbol{A}}_{t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$. By subtracting this identify from (B.2), we obtain
$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
@@
$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}=0,$ (B.4)
@@
Since $B_{y,t}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ by Lemma 1, we can solve $e_{t}$ as $e_{t}=C\tilde{e}_{t}$, where $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ is $\mathcal{F}_{t-1}$-measurable. By Assumption 1, the random variables $\tilde{e}_{1t},...,\tilde{e}_{dt}$ are mutually independent and at most one of them is Gaussian. Also the random variables $e_{1t},...,e_{dt}$ are mutually independent and at most one of them is Gaussian. Therefore, conditionally on $\mathcal{F}_{t-1}$, by Darmois-Skitovich theorem (Lemma A.1 in Lanne et al., 2017), at most on column of $C$ may contain more than one nonzero element. Suppose, say, the $k$th column of $C$ has at least two nonzero elements, $c_{ik}$ and $c_{jk}$, $i\neq j$. Then, $e_{it}=c_{ik}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{il}\tilde{e}_{lt}$ and $e_{jt}=c_{jk}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{jl}\tilde{e}_{lt}$ with $\tilde{e}_{kt}$ Gaussian (by Darmois-Skitovich theorem). This implies that $E[e_{it}e_{jt}]=c_{ik}c_{jk}\neq 0$ (since $E[\tilde{e}_{kt}^{2}]=1$) so that $e_{it}$ and $e_{jt}$ are not independent, which is a contradiction. Hence, each column of $C$ has at most one nonzero element.
@@
Threshold weights. To show the ”moreover”-part, consider first the threshold transition weights defined in (2.3). Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$. That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
@@
At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{0,m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$. Consequently, Equation (B.4) can be written as
$\tilde{\phi}_{\tilde{m}_{t},0}-\phi_{m_{t},0}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$.
@@
Now, Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t},0}-\phi_{m_{t},0}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$. Hence, by applying Lemma B.1 to each scalar equation in (B.6), it can be concluded that the set of solutions to Equation (B.6) has a Lebesgue measure zero in $\mathbb{R}^{dp}$ unless it is identically true. Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t},0}=\phi_{m_{t},0}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m},0}=\phi_{m,0}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m},0}=\phi_{m,0}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$. Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so that the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.

### v4

@@ 6 Conclusion
In addition to establishing the identification of the shocks, we have discussed the problem of labelling them, and employed blended identification (cf. Carriero et al., 2024) to address weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices. We have proposed estimating the model parameters by the method of penalized maximum likelihood, and introduced a three-step estimation procedure by adapting the approach of Lanne et al. (2017) to nonlinear SVAR models. Building on the results of Saikkonen (2008), we have also provided a sufficient condition for ergodic stationarity of our structural STVAR model. The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
In an empirical application to U.S. data from 1987:4 to 2024:12, we have studied the macroeconomic effects of the climate policy uncertainty shock using a two-regime logistic STVAR model. As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics. To distinguish between periods low and high economic policy uncertainty, we used the first lag of the EPU index (Baker et al., 2016) as the switching variable. Our results are in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024). However, we found that while a positive CPU shock decreases production and increases inflation in times of both low and high EPU, its inflationary effects are stronger during the periods of high EPU.
@@ Appendix A Stationarity and ergodicity
Establishing stationarity and ergodicity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties such as consistency of the estimator. Saikkonen (2008) derives a sufficient condition for ergodic stationarity of the smooth transition vector error correction (STVEC) model. Kheifets and Saikkonen (2020) make use of the results of Saikkonen (2008) to show that the stationarity condition readily applies for a STVAR model that can be obtained as a special case of the STVEC model of Saikkonen (2008). However, in both Saikkonen (2008) and Kheifets and Saikkonen (2020), the conditional covariance matrix of the process is defined as a weighted sum of positive definite error term covariance matrices of the regimes. As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model is not nested to the model of Saikkonen (2008). The applicability of the results of Saikkonen (2008) also requires that the weights are not exogenous but functions of the preceding $p$ observations (the time $t$ weights are allowed to depend on the time $t$ observation of an IID process independent of $e_{t}$, but we do not discuss this option further, as it violates our assumption of $\mathcal{F}_{t-1}$-measurable or nonrandom weights). Nonetheless, we will show in this section that the stationarity condition applies to our model as well.
@@
The following assumption collects the sufficient conditions for establishing ergodicity, stationarity, and mixing properties of our structural STVAR model.
@@
Condition (I) states that the JSR of the companion form AR matrices of the regimes is strictly less than one, and it is analogous to Condition (10) of Saikkonen (2008). Note that this condition is sufficient but not necessarily necessary for ergodic stationarity of the process (to the best of our knowledge, there does not currently exist results presenting a sufficient necessary condition for STVAR models). Condition (II) is analogous Assumption 1 of Saikkonen (2008), and it is innocuous in practice but rules out bounded error distributions, for example. Finally, Condition (III) ensures that the assumptions for the transition function imposed by Saikkonen (2008) are satisfied, ruling out nonrandom exogenous weights (which are accommodated by our identification results).
@@
Theorem A.1 is proven in Appendix A.1. A notable difference of Theorem A.1 to Theorem 1 of Saikkonen (2008) (and Theorem 1 of Kheifets and Saikkonen, 2020) is that our results holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, while the result of Saikkonen (2008) holds everywhere. To be more specific, the result of Theorem A.1 holds with such values of $B_{1},...,B_{M}$ that the resulting impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$, and Lemma 1 shows that this holds everywhere expect in regions of $\mathbb{R}^{d\times dM}$ that have a Lebesgue measure zero.
Assumption A.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption A.1(I) is computationally demanding to verify in practice (e.g., Chang and Blondel, 2013). This makes it poorly suitable for restricting the parameter space in the numerical estimation of parameters discussed in Section 4. During estimation, this condition would have to be verified repeatedly a very large of times, making the estimation extremely tedious. Hence, it is useful make use of the following condition in the estimation, which is necessary for Assumption A.1(I).
@@
Condition A.1 states that the usual stability condition is satisfied by each of the regimes. It is necessary for Assumption A.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein). After estimation, Assumption A.1(I) can be checked for the solutions of interest.
Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), but the implementation is computationally very demanding if the matrices are large and a tight bound is required. The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
@@ B.1 Proof of Lemma 1
Before proving Lemma 1, we establish the following lemma that is made use of in the proof:
@@
Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variables $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
@@
To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$. First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$. Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible. The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$. Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$. Denote this polynomial as $P_{t}(B_{1},...,B_{M})\equiv\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})$. Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero. It then follows from Lemma B.1 that $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
@@ B.2 Proof of Lemma 2
The first part of our proof is largely similar to the proof of Proposition 1 in Lanne et al. (2017) building on results from independent component analysis (Comon, 1994). First, note that from Equation (3.2), we can rewrite the model as
$y_{t}-\phi_{y,t}-\boldsymbol{A}_{y,t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (B.2)
where $\phi_{y,t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ $(d\times 1)$, $\boldsymbol{A}_{y,t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{y,t}-\tilde{\boldsymbol{A}}_{y,t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$. By subtracting this identify from (B.2), we obtain
$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
@@
$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=0,$ (B.4)
@@
Since $B_{y,t}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ by Lemma 1, we can solve $e_{t}$ as $e_{t}=C\tilde{e}_{t}$, where $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ is $\mathcal{F}_{t-1}$-measurable. By Assumption 1, the random variables $\tilde{e}_{1t},...,\tilde{e}_{dt}$ are mutually independent and at most one of them is Gaussian. Also the random variables $e_{1t},...,e_{dt}$ are mutually independent and at most one of them is Gaussian. Therefore, conditionally on $\mathcal{F}_{t-1}$, by Darmois-Skitovich theorem (Lemma A.1 in Lanne et al., 2017), at most one column of $C$ may contain more than one nonzero element. Suppose, say, the $k$th column of $C$ has at least two nonzero elements, $c_{ik}$ and $c_{jk}$, $i\neq j$. Then, $e_{it}=c_{ik}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{il}\tilde{e}_{lt}$ and $e_{jt}=c_{jk}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{jl}\tilde{e}_{lt}$ with $\tilde{e}_{kt}$ Gaussian (by Darmois-Skitovich theorem). This implies that $E[e_{it}e_{jt}]=c_{ik}c_{jk}\neq 0$ (since $E[\tilde{e}_{kt}^{2}]=1$), so $e_{it}$ and $e_{jt}$ are not independent, which is a contradiction. Hence, each column of $C$ has at most one nonzero element.
@@
Threshold weights. To show the ”moreover”-part, consider first the threshold transition weights defined in (2.3). Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$, $m=1,...,M$. By Condition (3) of Lemma 2, the sets $\mathcal{T}_{m}$ and $\tilde{\mathcal{T}}_{m}$ are nonempty for all $m=1,...,M$. That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
@@
At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$. Consequently, Equation (B.4) can be written as
$\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$ at the time $t$.
@@
Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$. Hence, by applying Lemma B.1 to each scalar equation in (B.6), it can be concluded that the set of solutions to Equation (B.6) has a Lebesgue measure zero in $\mathbb{R}^{dp}$ unless it is identically true. Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t}}=\phi_{m_{t}}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (as a countable union of measure zero sets has a measure zero). It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$. Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.